OpenStudy (3mar):

One has deposited money into a bank account that earned 1.25% simple interest each year. After 12 year, he had earned $5.00 in interest on the account. If no other money was deposited into or withdrawn from the account, how much was his initial deposit?

OpenStudy (3mar):

@mathmate If you could help, I would be grateful!

OpenStudy (queenbb):

i need heelp

OpenStudy (mathmate):

The formula for simple interest is I=Pin where I=acumulated interest P=principal i=annual interest (in decimal form) n=number of years. So putting numbers in, 5.0=P(0.0125)12 Solve for P.

OpenStudy (3mar):

Thank you for good explanation, @mathmate !! It really shows up it!

OpenStudy (3mar):

But it shows a decay and I think it is a growth! am I right, @mathmate?

OpenStudy (mathmate):

The question says 1. simple interest, so it is NOT an exponential growth (or decay), it means that interest for a given principal will accumulate the same interest year after year, there is no compounding, that is why I=Pin 2. The given amount is the interest, principal is not included, that is why the principal is much higher (6-7 times) than the interest over 12 months.

OpenStudy (mathmate):

The formula for combined principal and interest would therefore be: Accumulated amount=P+I=P+Pin=P(1+in)

OpenStudy (will.h):

We can use the formula A = P(1 + r)^x Since A is 5 and r is 1.25% and x is 12 years 5 = P(1 + 0.0125)^12 5 = p(1.0125)^12 5 = p(1.160754518) 4.307543002 = p So the initial investment is 4.307543002 The question is does it even worth to wait all that for such investment xD

OpenStudy (3mar):

"2. The given amount is the interest, principal is not included, that is why the principal is much higher (6-7 times) than the interest over 12 months."???? @mathmate

OpenStudy (mathmate):

Recall that we are working with 1.25% \(simple\) interest, even though this has disappeared from the market for years since computer age. FYI: http://www.investopedia.com/terms/s/simple_interest.asp P is principal, so solving for P, we have 5=P(0.0125)(12) P=5/(12/0.0125)=$33.33

OpenStudy (3mar):

Do you mean \[P=5/(12/0.0125)=$33.33\] or \[P=5/(12*0.0125)=$33.33\]??

OpenStudy (mathmate):

Sorry, the latter! I was rationalizing 0.0125 as 1/80 and dividing 12 by 80.

OpenStudy (3mar):

so is it * or / ?

OpenStudy (mathmate):

\(P=5/(12*0.0125)=$33.33\) is correct as the result of solving I=Pin => P=I/(in)

OpenStudy (3mar):

Oooooooh so the simple interest imply that it is just the amount of bulk × annual interest and that is for one year. If it is calculated for n years, it would be P×i×n I GOT IT! I am so much thankful to you!

OpenStudy (mathmate):

No problem! :)